The perimeter of a rectangle is 32cm. One side is 10 cm longer than the other side. Find the area of the rectangle.
Answer by fafa(13)

Let x1=length1 and x2=length2
x1=x1 x2=x1+10
perimeter = 2. (x1 +x2)
32 = 2.(x1 + (x1+10))
32 = 2.(2.x1 +10)
32 = 4.x1 + 20
32 - 20 = 4.x1 +20 -20
12 = 4.x1
x1= 12/4
x1 = 3
x2= x1+10=3+10=13
area= 3 . 13 = 39